OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, -1).

OpenStudy (waheguru):

Do you know how to solve this?

OpenStudy (anonymous):

nooo im soo lost

OpenStudy (waheguru):

What don't you understand

OpenStudy (anonymous):

i dont understand how to solve it. its been a long time since i did algebra and i cant remember how to do it

OpenStudy (waheguru):

Lets Do this step by step Do you know parallel and perpendicular lines ? And how the slope change in relation to each

OpenStudy (waheguru):

When a line is parallel to a line then the slope stays the same  Here as you can see the lines will never meet because the slopes are the same althought the y intercepts can be different, even if the slopes were 2 for the first line and 1.9 for the second they wound not be parallel *When a line is parallel to another line the slope is the same Perpendicular lines intersect at 90 degrees and for that to happend the slope must be the opposite reciprocal so if a line is parallel to 1/2, we find the reciprocal which is 2/1 and change it from a positive to negative so it becomes -2/1 Do you get this much

OpenStudy (waheguru):

Sorry I meant to say when a line is perpendicular for the second paragraph

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh okay

OpenStudy (waheguru):

3x + y = 7 Here you need to find the equation of the line which is perpendicular to this equation so right of the bat we need to make this in to y=mx+b format right so we can get the slope Do you know how to change it in to y=mx+b format?

OpenStudy (waheguru):

Convert 3x + y = 7 in to y=mx+b @drakelovermel

OpenStudy (anonymous):

i sorta forgot. please refresh my memory

OpenStudy (waheguru):

This is algebra 3x + y = 7 you need to isolate y amd move everythin g on the right side y=-3x+7

OpenStudy (waheguru):

Do you understand

OpenStudy (anonymous):

yes

OpenStudy (waheguru):

Now what is the perpendicular slope to -3x?

OpenStudy (anonymous):

was it y=3x+7

OpenStudy (waheguru):

No we moved it so it became negative

OpenStudy (anonymous):

ohh okay i got you

OpenStudy (waheguru):

So...

OpenStudy (anonymous):

okay so whats to learn next?

OpenStudy (waheguru):

What is the perpendicular slope to -3x.

OpenStudy (anonymous):

heyy thanks for the help! you where very helpful but i have to go eat and my mom is calling me soo im soo srry but i got to go. but thanks!!!

OpenStudy (waheguru):

Np!

OpenStudy (anonymous):

thank you soo much!